For the reaction shown here, 5.7 molA is mixed with 3.2 molB and 2.5 molC. What is the limiting reactant?

3A+2B+C→2D

A

B

C

D
Answer #1

3A+2B+C→2D

A: B : C = 3 : 2: 1

C is exceses reactant

B is 3.2 moles   = 3.2/2= 1.6 moles of B

5.7 moles of A = 5.7/3 = 1.9 moles of A

According to problem

A : B : C = 1.9 : 1.6: 2.5

Here limiting reagent is B